# Approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search

### v1

@@ 6 Conclusion
In this paper, we propose the approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search (ABPG-VMAW) for composite nonconvex optimization problems. Our line search condition allows a larger step-size than existing algorithms. We have established global subsequential convergence with some standard assumptions. We have guaranteed global convergence to a stationary point under the KL property even if $g\not\equiv 0$. This is the first contribution on ABPG-type algorithms. Moreover, our numerical experiments on $\ell_{p}$ regularized least squares problems and nonnegative linear inverse problems have shown that ABPG-VMAW outperforms ABPG and proximal gradient algorithms.
On the other hand, our line search procedure would not be well-defined when the objective function is dominated more by $g$ than by $f$ (in practice, this case is rare when $g$ is a regularizer). Although we establish that our line search is well-defined when $g\equiv 0$ in Section A, it would be important to prove that in the general case $g\not\equiv 0$.
@@ Appendix A Appendix: Implementation of Line Search
In order to obtain a step-size $t_{k}$ satisfying both (3.6) and (3.7), we adopt a bisection method for line search procedures in Algorithm 2.
@@ A.1 Special Case: $g\equiv 0$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$
From the continuity of $A(t)$, there exists a positive number $\Delta$ such that $A(t)<0$ holds for any $t$ in $t^{*},t^{*}+\Delta]$, and from the definition of $t^{*}$ there exists nonempty subset $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$ of $t^{*},t^{*}+\delta]$ such that $W(t)>0$ always holds i.e., (A.1) and (A.2) hold simultaneously on $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$. ∎
Lemma A.2 ensures that when the sign of $A(t)$ changes from negative to positive at two points, an Armijo–Wolfe step size exists between those two points.
@@
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. Whenever the second loop of the line search procedure in each iteration terminates, the final trial step $t$ is an Armijo–Wolfe step if $\lambda$ is small enough. If, on the other hand, the line search procedure does not terminate, then it eventually generates a nested sequence of finite intervals $\alpha,\beta]$, halving in length at each iteration, and each containing a set of nonzero measure of Armijo–Wolfe steps. These intervals converge to a step $t_{0}>0$ such that
@@
When $g\not\equiv 0$, Algorithm 2 would not terminate in finite steps to obtain $t_{k}$ such that (3.6) and (3.7) holds. For example, the influence of $g$ plays a principal role in determining the overall behavior of the objective function. However, this case is rare in practice because $g$ is often a regularizer. In fact, Algorithm 2 succeeds to obtain $t_{k}$ satisfying (3.6) and (3.7) in Section 5.2.

### v2

@@ 6 Conclusion
In this paper, we propose the approximate Bregman proximal gradient algorithm with variable metric Armijo–Wolfe line search (ABPG-VMAW) for composite nonconvex optimization problems. Our line search condition allows a larger step-size than existing algorithms. We have established global subsequential convergence under standard assumptions. Moreover, under the KL property, we have proved global convergence to a stationary point even when $g\not\equiv 0$. To the best of our knowledge, this is the first global convergence result for ABPG-type algorithms in this setting. Moreover, our numerical experiments on $\ell_{p}$ regularized least squares problems, nonnegative linear inverse problems, and phase retrieval have shown that ABPG-VMAW outperforms ABPG and proximal gradient algorithms.
On the other hand, our line search procedure would not be well-defined when the objective function is dominated by $g$ rather than by $f$ (in practice, this case is rare when $g$ is a regularizer). Although we establish that our line search is well-defined when $g\equiv 0$ in Section A, it is important to prove this in the general case $g\not\equiv 0$.
@@ Appendix A Appendix: Implementation of Line Search
In order to obtain a step-size $t_{k}$ satisfying both (3.6) and (3.7), we adopt a bisection method for the line search procedure in Algorithm 2.
@@ A.1 Special Case: $g\equiv 0$ and $\operatorname{dom}\phi=\mathbb{R}^{n}$
From the continuity of $A(t)$, there exists a positive number $\Delta$ such that $A(t)<0$ holds for any $t$ in $t^{*},t^{*}+\Delta]$, and from the definition of $t^{*}$ there exists a nonempty subset $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$ of $t^{*},t^{*}+\delta]$ such that $W(t)>0$ always holds i.e., (A.1) and (A.2) hold simultaneously on $\tilde{t}_{\alpha},\tilde{t}_{\beta}]$. ∎
Lemma A.2 ensures that when the sign of $A(t)$ changes from negative to positive at two points, an Armijo–Wolfe step-size exists between those two points.
@@
Suppose that Assumptions 3.1, 3.2, and 4.1 hold. Whenever the second loop of the line search procedure in an iteration terminates, the final trial step $t$ is an Armijo–Wolfe step, provided that $\lambda$ is small enough. If, on the other hand, the line search procedure does not terminate, then it eventually generates a nested sequence of finite intervals $\alpha,\beta]$, halving in length at each iteration, and each containing a set of nonzero measure of Armijo–Wolfe steps. These intervals converge to a step $t_{0}>0$ such that
@@
When $g\not\equiv 0$, Algorithm 2 would not terminate in finite steps to obtain $t_{k}$ such that (3.6) and (3.7) hold. For example, the influence of $g$ plays a principal role in determining the overall behavior of the objective function. However, this case is rare in practice because $g$ is often a regularizer. In fact, Algorithm 2 succeeds in obtaining $t_{k}$ satisfying (3.6) and (3.7) in Section 5.2.
Remark A.5.
Let $x\in\operatorname{int}\operatorname{dom}\phi$, $d\in\mathbb{R}^{n}$, and $t\in\mathbb{R}_{+}$. Suppose $d\neq 0$ because $\langle\nabla f(x+td)+\xi,d\rangle\geq c_{2}\langle\nabla f(x)+\xi,d\rangle$ always holds when $d=0$. We assume that $\phi$ is $\sigma$-strongly convex (see Assumption 4.1) and $g$ is $\kappa$-Lipschitz continuous with $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$, i.e., $\max_{\xi\in\partial g(x)}\|\xi\|\leq\kappa$ for any $x\in\operatorname{dom}g$. When $g$ is strictly continuous on $\operatorname{dom}g$ (e.g., locally Lipschitz continuous functions are strictly continuous), its subdifferential at any point of $\operatorname{dom}g$ is bounded [rockafellar2009variational, Theorem 9.13]. Using this and letting $\xi\in\partial g(x)$ and $\xi^{+}\in\partial g(x+td)$, we have
$0\leq\langle\xi^{+}-\xi,td\rangle\leq(\|\xi^{+}\|+\|\xi\|)\|td\|\leq 2\kappa t\|d\|,$ (A.15)
which implies
$\langle\xi^{+}-\xi,d\rangle\leq 2\kappa\|d\|\leq\frac{c_{1}\sigma}{2\lambda}\|d\|^{2}\leq\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle,$ (A.16)
where the second inequality holds from $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$, and the last inequality holds because of the strong convexity of $\phi$.
Suppose that $t$ satisfies the following Armijo condition:
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $\geq c_{1}\left(\langle\nabla f(x),d\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right).$ (A.17)
The existence of such $t$ is assured by a well-known discussion, e.g.[Nocedal2006-sb], and the mean-value theorem. We obtain
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $\geq c_{1}\left(\langle\nabla f(x),d\rangle+g(x+d)-g(x)+\frac{1}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle\right)$ (A.18)
$\geq c_{1}\langle\nabla f(x)+\xi,d\rangle+\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle,$ (A.19)
where the last inequality holds from the convexity of $g$. On the other hand, using (A.16), we also have
$\langle\nabla f(x+td)+\xi^{+},d\rangle$ $=\langle\nabla f(x+td)+\xi,d\rangle+\langle\xi^{+}-\xi,d\rangle$ (A.20)
$\leq\langle\nabla f(x+td)+\xi,d\rangle+\frac{c_{1}}{2\lambda}\langle\nabla^{2}\phi(x)d,d\rangle.$ (A.21)
Combining these inequalities and using $\langle\nabla f(x)+\xi,d\rangle<0$ from (4.2) and $c_{1}<c_{2}$ implies
$\langle\nabla f(x+td)+\xi,d\rangle\geq c_{1}\langle\nabla f(x)+\xi,d\rangle\geq c_{2}\langle\nabla f(x)+\xi,d\rangle.$ (A.22)
Therefore, our line search procedure is well-defined. However, $\kappa\leq\frac{c_{1}\sigma}{4\lambda}\|d\|$ does not holds when $g$ is a dominance term, i.e., $\kappa>\frac{c_{1}\sigma}{4\lambda}\|d\|$ holds. In this case, our line search procedure might not be well-defined.